OpenStudy (anonymous):

y=sec(x/2-3pi/4) + 1 Period: Phase Shift: Domain: Range: Asymptotes: Graph

OpenStudy (shamil98):

hold on, F(t) = Asin(Bt – C) + D There is no amplitude in this.. The Period is 2PI/B 2pi/x/2 = 2pi(2/x) = 4pi/x - C = phase shift - (-3pi/4) = 3pi/4 units That's all i remember..

OpenStudy (anonymous):

Thanks!